The salesclerk at a clothing store earns a 16% commission on sales. If the clerk sold 24 purses at $60 each, how much commission did he earn?

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52814)

.

Total sales is  24*60 = 1440  dollars.

The commission he will earn is 16% of it, i.e.  0.16*1440 = 230.40 dollars.    ANSWER

Solved.



Answer by Theo(13342)

commission is 16% of sales.
24 purses * 60 = total sales of 1440
.16 * 1440 = 230.4

clerk made 230.4 in commission.